# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

### v1

@@ E.1 Pre-training Datasets
Real-world data. The real-world datasets were stratified into five tiers based on their predictability. This hierarchical structure dictates the sampling probability during model training, assigning a higher likelihood of selection to datasets with greater predictability. Such a strategy ensures that the model is preferentially trained on high-quality data while preserving its capacity to predict corner cases. Specifically, Tier 1 comprises datasets characterized by pronounced periodicity and trends with low noise. Tier 2 contains datasets with similarly distinct patterns but high noise, whereas Tier 3 includes those with subtle trends and considerable noise. The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity. The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 5-8.
@@
Synthetic data. We build a synthetic data generator that produces two distinct types of time series, each with a length of 4096. The first type consists of composite series, created by the additive combination of seasonal, trend, and noise components. For seasonality, we sample one or two components, where the primary period is drawn from {24, 48, 288, 360}, and a potential second period is a fixed seven-fold multiple of the first; the seasonal patterns manifest as either spike trains or smooth non-sinusoidal templates generated via interpolation. An optional trend is chosen from linear, exponential, or an ARIMA-like process derived from cumulating a stationary ARMA model’s output. High-probability white Gaussian noise is also added, with each series guaranteed to contain at least one seasonal or trend component and all magnitudes bounded for stability. The generation process for these composite series is formally detailed in Algorithm 2. Complementing these are idealized industrial signals, which simulate perfectly regular machine cycles. These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3.
@@
SZ-Taxi Transport 15T 156 464,256
M DENSE Transport 15T 30 525,600
Electricity (15 Min.) Energy 15T 370 41,936,458
@@
COVID Deaths Healthcare D 266 56,392
Temperature Rain Nature D 32,072 22,290,040
@@
1:Time series length $L=4096$, Primary period set $\mathcal{P}_{1}=\{24,48,288,360\}$, Harmonic multiplier $n=7$, Trend types $\mathcal{T}_{\text{types}}=\{\text{linear, exp, ARMA}\}$, Seasonal patterns $\mathcal{S}_{\text{patterns}}=\{\text{spike, interpolated segment}\}$.
2:A synthetic time series $\bm{x}_{1:L}$.
@@
4:Sample flags $f_{s},f_{t},f_{n}$ $\triangleright$ Determine component inclusion, ensuring $f_{s}\lor f_{t}$ is true
5:if $f_{s}$ is true then
6:  $k\sim\text{Bernoulli}(0.2)$ $\triangleright$ $k=1$ for double period (20% prob), $k=0$ for single
7:  $p_{1}\sim\mathcal{U}(\mathcal{P}_{1})$ $\triangleright$ Sample primary period
8:  $\mathcal{P}_{\text{active}}\leftarrow\{p_{1}\}$
9:  if $k=1$ then
10:   $\mathcal{P}_{\text{active}}\leftarrow\mathcal{P}_{\text{active}}\cup\{n\cdot p_{1}\}$
11:  end if
12:  for $p$ in $\mathcal{P}_{\text{active}}$ do
13:   $a\sim\mathcal{U}(1.0,3.0)$ $\triangleright$ Sample amplitude for this component
14:   $pattern\sim\mathcal{U}(\mathcal{S}_{\text{patterns}})$
15:   $\bm{c}\leftarrow\text{GeneratePattern}(pattern,p,a)$ $\triangleright$ Create a single cycle of the pattern
16:   $\bm{s}\leftarrow\bm{s}+\text{Tile}(\bm{c},L)$ $\triangleright$ Tile the cycle to length $L$ and add to seasonal component
17:  end for
18:end if
19:if $f_{t}$ is true then
20:  $type\sim\mathcal{U}(\mathcal{T}_{\text{types}})$
21:  $\bm{t}\leftarrow\text{GenerateTrend}(type,L)$
22:  if $f_{s}$ is true then
23:   $\lambda\sim\mathcal{U}(0.1,0.3)$ $\triangleright$ Reduce trend strength when seasonality is present
24:   $\bm{t}\leftarrow\lambda\cdot\bm{t}$
25:  end if
26:end if
27:if $f_{n}$ is true then
28:  $\sigma_{n}\sim\mathcal{U}(0.01,0.1)$
29:  $\bm{n}\sim\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Generate white Gaussian noise
30:end if
@@
1:Time series length $L=4096$, Pattern types $\mathcal{P}_{\text{types}}=\{\text{inverted\_u, spikes}\}$.
2:A synthetic time series $\bm{x}_{1:L}$.
@@
7:if $type=\text{inverted\_u}$ then
8:  $sign\leftarrow-1$
9:else
10:  $sign\leftarrow+1$
11:end if
12:for $i\leftarrow 0,p,2p,\dots$ up to L1L-1 do
13:  $start\leftarrow i$, $end\leftarrow\min(i+w,L)$
14:  $\bm{x}_{start:end}\leftarrow\bm{x}_{start:end}+sign\cdot\bm{e}_{1:end-start}$ $\triangleright$ Add or subtract event shape at periodic intervals
15:end for
16:if $\sigma_{n}>0$ then
17:  $\bm{x}\leftarrow\bm{x}+\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Add global Gaussian noise
18:end if
19:return $\bm{x}_{1:L}$
@@ E.3 Evaluation Length Selection
E.4 Details of IARoPE Analysis
In this section, we explain in more detail the setting of the IARoPE analysis experiment discussed in Section 4.4.2. To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of “shuffle” experiments. These experiments function as a form of causal intervention analysis. The underlying hypothesis is: if the instance-derived $\theta$ modulations are crucial for IARoPE’s performance, then disrupting this linkage by applying $\theta$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
@@ E.4 Details of IARoPE Analysis
  • IARoPE: Our proposed method, where each time series instance utilizes its own FFT-derived features to independently modulate its RoPE $\theta$ parameters at each layer.
  • Intra-Dataset Shuffle: During inference, the learned layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly permuted among instances within the same batch and originating from the same dataset. The modulated $\theta$ are always shuffled in the same layer.
  • Inter-Dataset Shuffle: For instances of a target dataset, layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly sampled from a pre-computed collection derived from a different source dataset and applied to instances of the target dataset. This process also ensures layer-wise correspondence of the applied modulations.
  • Fixed RoPE: We use the fully trained IARoPE model but disable its instance-adaptive modulation during inference. Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta_{\text{init}}$ across all layers. This approach keeps the base model architecture and other learned weights identical to IARoPE, isolating the effect of the modulation.
By comparing IARoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta$ modulation process.
@@ Appendix F Full Evaluation Results
$\textsc{Kairos}_{b}$ MSE 0.396 0.434 0.452 0.426 0.427 0.276 0.353 0.385 0.386 0.350 0.295 0.336 0.367 0.393 0.348 0.160 0.220 0.271 0.357 0.252 0.146 0.192 0.248 0.338 0.231 0.927 0.955 0.945 0.989 0.954 0.817 0.843 0.804 0.821 0.471 28
MAE 0.385 0.408 0.418 0.427 0.410 0.319 0.369 0.394 0.412 0.374 0.322 0.355 0.380 0.402 0.365 0.238 0.283 0.317 0.373 0.303 0.182 0.228 0.270 0.330 0.253 0.414 0.425 0.427 0.432 0.425 0.446 0.449 0.434 0.443 0.364
Appendix G Showcases
@@
Appendix H Broader Impacts
Our work on Kairos is foundational research focused on advancing time series modeling. While Kairos itself is not designed for direct societal applications with immediate negative impacts, any powerful predictive technology could be misused. So beyond general risks of advanced AI, our model has no specific negative societal impacts need to be discussed here.

### v2

@@ E.1 Pre-training Datasets
Real-world data. The real-world datasets were stratified into five tiers based on their predictability. This hierarchical structure dictates the sampling probability during model training, assigning a higher likelihood of selection to datasets with greater predictability. Such a strategy ensures that the model is preferentially trained on high-quality data while preserving its capacity to predict corner cases. Specifically, Tier 1 comprises datasets characterized by pronounced periodicity and trends with low noise. Tier 2 contains datasets with similarly distinct patterns but high noise, whereas Tier 3 includes those with subtle trends and considerable noise. The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity. The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 6-9.
@@
Synthetic data. We build a synthetic data generator that produces two distinct types of time series, each with a length of 4096. The first type consists of composite series, created by the additive combination of seasonal, trend, and noise components. For seasonality, we sample one or two components, where the primary period is drawn from {24, 48, 288, 360}, and a potential second period is a fixed seven-fold multiple of the first; the seasonal patterns manifest as either spike trains or smooth non-sinusoidal templates generated via interpolation. An optional trend is chosen from linear, exponential, or an ARIMA-like process derived from cumulating a stationary ARMA model’s output. High-probability white Gaussian noise is also added, with each series guaranteed to contain at least one seasonal or trend component and all magnitudes bounded for stability. The generation process for these composite series is formally detailed in Algorithm 2. Complementing these are idealized industrial signals, which simulate perfectly regular machine cycles. These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3. Synthetic dataset cases are provided in Appendix H.1.
Distribution Analysis. To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto (Cohen et al., 2025). We employed the ARCH-LM Statistic and Spectral Entropy to quantify time-varying volatility and information density (complexity), respectively. These metrics align closely with the design motivation for the Mixture-of-Size Encoding and DRoPE components of Kairos. As illustrated in Figures 6 (a) and (b), we observed that a significant portion of the real-world dataset exhibits ARCH-LM Statistic values approaching zero and Spectral Entropy values nearing one, indicating high time-varying volatility and complexity. In contrast, the synthetic dataset exhibits the opposite characteristics, exposing the model during training to time series with more constant volatility and greater regularity. This exposure enhances the model’s generalization capabilities. Furthermore, we statistically compared the periodicity of the real-world and synthetic datasets by plotting their respective Cumulative Distribution Functions (CDFs). As shown in Figure 6 (c), the dominant periods in the real-world dataset are heavily concentrated. This can cause the model to overfit to these specific periodicities and fail to learn generalizable periodic patterns. Conversely, the synthetic dataset presents a significantly smoother distribution of periods, enabling the model to generalize effectively across arbitrary periodicities. Consequently, the synthetic dataset effectively mitigates these biases present in the real-world data, thereby exposing the model to a more comprehensive and diverse set of distributions.
Figure 6: Comparison of real and synthetic datasets across (a) ARCH-LM statistic, (b) spectral entropy, and (c) dominant period distributions, illustrating complementary volatility, complexity, and periodicity characteristics.
@@
0: Time series length $L=4096$, Primary period set $\mathcal{P}_{1}=\{24,48,288,360\}$, Harmonic multiplier $n=7$, Trend types $\mathcal{T}_{\text{types}}=\{\text{linear, exp, ARMA}\}$, Seasonal patterns $\mathcal{S}_{\text{patterns}}=\{\text{spike, interpolated segment}\}$.
0: A synthetic time series $\boldsymbol{x}_{1:L}$.
@@
2: Sample flags $f_{s},f_{t},f_{n}$ {Determine component inclusion, ensuring $f_{s}\lor f_{t}$ is true}
3:if $f_{s}$ is true then
4:  $k\sim\text{Bernoulli}(0.2)$ {$k=1$ for double period (20% prob), $k=0$ for single}
5:  $p_{1}\sim\mathcal{U}(\mathcal{P}_{1})$ {Sample primary period}
6:  $\mathcal{P}_{\text{active}}\leftarrow\{p_{1}\}$
7:  if $k=1$ then
8:   $\mathcal{P}_{\text{active}}\leftarrow\mathcal{P}_{\text{active}}\cup\{n\cdot p_{1}\}$
9:  end if
10:  for all $p$ in $\mathcal{P}_{\text{active}}$ do
11:   $a\sim\mathcal{U}(1.0,3.0)$ {Sample amplitude for this component}
12:   $pattern\sim\mathcal{U}(\mathcal{S}_{\text{patterns}})$
13:   $\boldsymbol{c}\leftarrow\text{GeneratePattern}(pattern,p,a)$ {Create a single cycle of the pattern}
14:   $\boldsymbol{s}\leftarrow\boldsymbol{s}+\text{Tile}(\boldsymbol{c},L)$ {Tile the cycle to length $L$ and add to seasonal component}
15:  end for
16:end if
17:if $f_{t}$ is true then
18:  $type\sim\mathcal{U}(\mathcal{T}_{\text{types}})$
19:  $\boldsymbol{t}\leftarrow\text{GenerateTrend}(type,L)$
20:  if $f_{s}$ is true then
21:   $\lambda\sim\mathcal{U}(0.1,0.3)$ {Reduce trend strength when seasonality is present}
22:   $\boldsymbol{t}\leftarrow\lambda\cdot\boldsymbol{t}$
23:  end if
24:end if
25:if $f_{n}$ is true then
26:  $\sigma_{n}\sim\mathcal{U}(0.01,0.1)$
27:  $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma_{n}^{2}\mathbf{I})$ {Generate white Gaussian noise}
28:end if
@@
0: Time series length $L=4096$, Pattern types $\mathcal{P}_{\text{types}}=\{\text{inverted\_u, spikes}\}$.
0: A synthetic time series $\boldsymbol{x}_{1:L}$.
@@
5:if $type=\text{inverted\_u}$ then
6:  $sign\leftarrow-1$
7:else
8:  $sign\leftarrow+1$
9:end if
10:for all $i\leftarrow 0,p,2p,\dots$ up to L1L-1 do
11:  $start\leftarrow i$, $end\leftarrow\min(i+w,L)$
12:  $\boldsymbol{x}_{start:end}\leftarrow\boldsymbol{x}_{start:end}+sign\cdot\boldsymbol{e}_{1:end-start}$ {Add or subtract event shape at periodic intervals}
13:end for
14:if $\sigma_{n}>0$ then
15:  $\boldsymbol{x}\leftarrow\boldsymbol{x}+\mathcal{N}(\boldsymbol{0},\sigma_{n}^{2}\mathbf{I})$ {Add global Gaussian noise}
16:end if
17:return $\boldsymbol{x}_{1:L}$
@@ E.3 Evaluation Length Selection
E.4 Details of DRoPE Analysis
In this section, we explain in more detail the setting of the DRoPE analysis experiment discussed in Section 4.4.2. To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of ?shuffle? experiments. These experiments function as a form of causal intervention analysis. The underlying hypothesis is: if the instance-derived $\theta_{\text{inst}}$ modulations are crucial for DRoPE’s performance, then disrupting this linkage by applying $\theta_{\text{inst}}$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
@@ E.4 Details of DRoPE Analysis
  • DRoPE: Our proposed method, where each time series instance utilizes its own FFT-derived features to independently modulate its RoPE $\theta_{\text{inst}}$ parameters at each layer.
  • Intra-Dataset Shuffle: During inference, the learned layer-specific $\theta_{\text{inst}}$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly permuted among instances within the same batch and originating from the same dataset. The modulated $\theta_{\text{inst}}$ are always shuffled in the same layer.
  • Inter-Dataset Shuffle: For instances of a target dataset, layer-specific $\theta_{\text{inst}}$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly sampled from a pre-computed collection derived from a different source dataset and applied to instances of the target dataset. This process also ensures layer-wise correspondence of the applied modulations.
  • Fixed RoPE: We use the fully trained DRoPE model but disable its instance-adaptive modulation during inference. Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta$ across all layers. This approach keeps the base model architecture and other learned weights identical to DRoPE, isolating the effect of the modulation.
By comparing DRoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
@@ Appendix F Full Evaluation Results
$\textsc{Kairos}_{b}$ MSE 0.396 0.446 0.472 0.434 0.437 0.270 0.348 0.367 0.376 0.340 0.291 0.337 0.368 0.408 0.351 0.163 0.224 0.274 0.361 0.255 0.147 0.191 0.248 0.333 0.230 0.936 0.972 0.963 1.014 0.971 0.786 0.817 0.766 0.790 0.471 26
MAE 0.385 0.414 0.429 0.432 0.415 0.325 0.378 0.395 0.410 0.377 0.321 0.355 0.378 0.405 0.365 0.239 0.283 0.317 0.376 0.304 0.181 0.227 0.270 0.326 0.251 0.406 0.424 0.428 0.437 0.424 0.445 0.446 0.429 0.440 0.365
Appendix G Calculation of information density
$H_{SE}(\mathbf{w}_{i})=-\sum_{k=0}^{M-1}p_{i}[k]\log_{2}(p_{i}[k]).$ (34)
$\mu_{SE}=\frac{1}{N}\sum_{i=1}^{N}H_{SE}(\mathbf{w}_{i}).$ (35)
Appendix H Showcases
H.1 Showcases of Synthetic Data
Figure 7 showcases representative examples generated by the algorithm using the synthetic dataset. As detailed in Appendix E.1, the generated synthetic dataset is classified into two categories. The first is a composite type, formed from a combination of seasonal, trend, and noise components, which is designated Custom in the figures. The second consists of idealized industrial signals, designated Perfect periodic.
Figure 7: Several distinct case types are generated by the synthetic dataset algorithm. Specifically, the Custom designation refers to a composite signal type, which is systematically constructed by combining seasonal, trend, and noise components. Concurrently, the Perfect periodic designation denotes synthesized, idealized industrial signals.
H.2 Showcases of Kairos